A distribution of values is normal with a mean of 180.9 and a standard deviation of 83.6.
Find P42, which is the score separating the bottom 42% from the top 58%.
P42 =

Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.

Answer by Theo(13342)

mean is 180.9
standard deviation is 83.6
z-score with 42% of the area under the normal distribution curve to the left of it is equal to -.2018934725.
raw score associated with that is obtained from the z-score formula.
z = (x - m) / s becomes -.2018934725 = (x - 180.9) / 83.6.
solve for x to get x = 164.0217057.
here's what it looks like on a graph.